OpenStudy (anonymous):

In a survey it was found that 21 people liked product A, 26 liked product B and 29 liked product C. If 14 people liked products A and B, 12 people liked products C and A, 14 people liked products B and C and 8 liked all the three products. Find how many liked product C only.

OpenStudy (amistre64):

venns are good for this, not required, but good nonetheless. start with the middle and work your way out

OpenStudy (amistre64):

In a survey it was found that: 21 A 26 B 29 C 14 AnB 12 AnC 14 BnC 8 all the three products subtract 8 from each of the "ands" 14 - 8: AnB = 6 12 - 8: AnC = 4 14 - 8: BnC = 6 6 8 4 6 subtract the "3 numbers" from each "single" set 21 - (14+4): A = 3 26 - (14+6): B = 6 29 - (14+4): C = 11

OpenStudy (amistre64):